OpenStudy (austinl):

Two bacteria are placed in a petri dish with unlimited resources and grow unchecked by the law of exponential growth. If they divide once every 20 minutes, how many bacteria are present after one day? Write an equation that solves this problem and use "Maple" to find the numerical answer. NOTE: Maple is a mathematics program, I was just copying the question verbatim. Disregard that portion of the question if you are unfamiliar with the program!

OpenStudy (austinl):

On another note, this is very likely a differential equations problem.

OpenStudy (amistre64):

this one?

OpenStudy (austinl):

Indeed.

OpenStudy (amistre64):

Two bacteria are placed in a petri dish with unlimited resources and grow unchecked by the law of exponential growth. If they divide once every 20 minutes, how many bacteria are present after one day? t b 0 2 20 4 <---- 24 hours is in here 40 8 80 16

OpenStudy (amistre64):

\[2P = Pe^{20r}\] \[2 = e^{20r}\] \[ln(2) = 20r\] \[\frac{ln(2)}{20} = r\] \[\Large A=2e^{t~ln(2)/20}\]

OpenStudy (amistre64):

3*24 = 60+12 = 72 periods of 20 minutes in a 24hr period

OpenStudy (amistre64):

i spose the t i sin minutes here ....

OpenStudy (austinl):

Ok, so \(t\) is the number of times that 20 minutes passes? And \(A\) is the number of bacteria present after one full day of the 20 minute cycles?

OpenStudy (amistre64):

if we use the equation:\[\Large A=2e^{\frac{ln2}{20}t}\] then at t=20, we have the 4 at t=40, we have the 8 etc ... so my random thought about t was askew. let t be the number of minutes that have passed

OpenStudy (amistre64):

there are 24*60 minutes in a day, so let t=24*60

OpenStudy (austinl):

Ok, very cool. So now all I need to do is figure out how the heck to input this into the program to solve it... and I realize that It would be easy to solve, but I am required to use the program. It is part of my grade, and a massive waste of time as well I might add. Thanks very much!

OpenStudy (amistre64):

yeah, good luck with the mapling ;)

OpenStudy (austinl):

Yeah, thanks lol!